OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w

OpenStudy (zzr0ck3r):

subtract 2l from both sides then tell me what you have

OpenStudy (anonymous):

I think so im really not good in math.

OpenStudy (anonymous):

you can do this. first, subtract 2l from both sides of the equation

OpenStudy (zzr0ck3r):

im telling you to do something. subtract 2l from both sides of your equation, and then show us what you have.

OpenStudy (anonymous):

lol, go for it zz

OpenStudy (anonymous):

Is it 0?

OpenStudy (zzr0ck3r):

is what 0?

OpenStudy (jhannybean):

\(\large P-2l=-2l+2l+2w\)

OpenStudy (jhannybean):

simplify.

OpenStudy (jhannybean):

what do you get? whats the right side turn into?

OpenStudy (anonymous):

6w?

OpenStudy (jhannybean):

what is -2 + 2?

OpenStudy (anonymous):

But these are the answer choices A. W=P-2l B. w=p/2l C. w= 2p/21 D. W= p-21/2

OpenStudy (anonymous):

and -2 + 2? is 0

OpenStudy (jhannybean):

We're going step by step. Not jumping to answers.

OpenStudy (anonymous):

-2 + 2? is 0

OpenStudy (luigi0210):

Good luck you two!

OpenStudy (jhannybean):

Yes. So you have \(\large P-2l=2w\).Now divide both sides of the equation by 2. \[\large\frac{P-2l}{2}=\frac{2w}{2}\]\[\large w=\frac{P-2l}{2}\]

OpenStudy (anonymous):

Ummm is it -1p?

OpenStudy (jhannybean):

nope.How did you get that? Explain?

OpenStudy (anonymous):

I divided -2 by 2

OpenStudy (jhannybean):

why?

OpenStudy (jhannybean):

Oh like that.Compare what we have accomplished to your answer choices.

OpenStudy (jhannybean):

I don't think you need to simplify any more.

OpenStudy (anonymous):

Because of this \[\frac{ p-2 }{ 2 }\]

OpenStudy (jhannybean):

But these are the answer choices A. W=P-2l B. w=p/2l C. w= 2p/21 D. W= p-21/2

OpenStudy (jhannybean):

So which answer choice do you think fits what we solved for?